the area of a triangle is 152 m squared, and the height is 16 m. what us the base?
Answer by nerdybill(7384)

the area of a triangle is 152 m squared, and the height is 16 m. what us the base?
.
Since area of any triangle is
(1/2)bh
we get:
152 = (1/2)bh
152 = (1/2)b(16)
152 = b(8)
19 m = b (base)